# Grav. Cosmol. No. issue, year

### v1

@@ 2.2 Observational data on the temperatures of the Earth and Mars
The second, more important item is the effect of salts dissolved in the martian water on its freezing point. If the water would be fresh (unsalted), then its freezing point should equal $0\mbox{${}^{\circ}$C}$ or 273 K. However, it is well known that martian water can be strongly salted, i.e., represent a brine [30]. Therefore, its freezing point might be substantially reduced. The freezing temperature of terrestrial oceans can be reduced down to approximately $2\mbox{${}^{\circ}$C}$ [32]; but the ancient martian ocean might be much more salted and, therefore, its freezing temperature reduced much stronger. For example, the freezing point can be as low as $21\mbox{${}^{\circ}$C}$ for NaCl brine and $55\mbox{${}^{\circ}$C}$ for CaCl2 brine at the optimal concentrations [33]. Unfortunately, we do not have any information about a chemical composition of the ancient martian ocean and, therefore, can only consider a set of various options. In Fig. 1, a series of different freezing temperatures—from $0\mbox{${}^{\circ}$C}$ to $50\mbox{${}^{\circ}$C}$—is shown by the violet bars. The upward arrows denote that a liquid ocean could exist above the corresponding temperatures. (Let us emphasize that temperatures of the Earth and Mars in this figure are normalized to the different values: $T_{\rm Ms0}=218\,K=-55\mbox{${}^{\circ}$}$ and $T_{\rm Es0}=288\,K=15\mbox{${}^{\circ}$C}$, respectively [23].)
@@ 2.3 Analysis for the entire lifetime of the Earth and Mars
While the upper left panel in Fig. 1 represents a summary of observational data, three other panels demonstrate a confrontation of this data with theoretical curves (8) for three different cosmological models, characterized by the relative densities of the ‘dark energy’ ${\Omega}_{\Lambda}$ and non-relativistic (dust-like) matter ${\Omega}_{\rm D}$, both ‘dark’ and visible. (The Universe is always assumed to be spatially flat, so that ${\Omega}_{\Lambda}+{\Omega}_{\rm D}=1$.) The theoretical dependences $T_{\rm s}(t)/T_{\rm s0}$ are drawn by the black dotted curves, and the corresponding values of the contemporary Hubble parameter $H_{0}$ are written near each curve. For concineses, these values are normalized to $h=100$ km/s/Mpc.
In the case of commonly-accepted cosmological model with ${\Omega}_{\Lambda}=0.7$ and ${\Omega}_{\rm D}=0.3$ (bottom left panel), the best fit is achieved at $H_{0}=1.2{-}1.5~h=120{-}150$ km/s/Mpc, i.e., almost two times greater than the standard value of Hubble parameter at the intergalactic scales. The martian data (violet bars) suggest that such values of $H_{0}$ are consistent with the assumption of weakly-salted ocean, whose freezing point was above $10\mbox{${}^{\circ}$C}$. Significant outliers in the terrestrial data at the late times (within 1 billion years from today) are not surprising: this period involved an active biological evolution, resulting in the considerable variations of coefficients $A$ and $\alpha$ in our model (6, which were not taken into account.
@@
On the other hand, in the opposite case when the entire cosmological background is formed by the non-relativistic (dust-like) matter (${\Omega}_{\Lambda}=0$ and ${\Omega}_{\rm D}=1$, bottom right panel), the best-fitting Hubble parameter can be somewhat reduced: $H_{0}=1.0{-}1.2~h=100{-}120$ km/s/Mpc. Unfortunately, the corresponding theoretical curves become noticeably concaved and fit the experimental data worse at the later times.

### v2

@@ 2.2 Observational data on the temperatures of the Earth and Mars
The second, more important item is the effect of salts dissolved in the martian water on its freezing point. If the water would be fresh (unsalted), then its freezing point should equal $0\mbox{${}^{\circ}$C}$ or 273 K. However, it is well known that martian water can be strongly salted, i.e., represent a brine [30]. Therefore, its freezing point might be substantially reduced. The freezing temperature of terrestrial oceans can be reduced down to approximately $2\mbox{${}^{\circ}$C}$ [32]; but the ancient martian ocean might be much more salted and, therefore, its freezing temperature reduced much stronger. For example, the freezing point can be as low as $21\mbox{${}^{\circ}$C}$ for NaCl brine and $55\mbox{${}^{\circ}$C}$ for CaCl2 brine at the optimal concentrations [33]. Unfortunately, we do not have any information about a chemical composition of the ancient martian ocean and, therefore, can only consider a set of various options. In Fig. 1, a series of different freezing temperatures—from $0\mbox{${}^{\circ}$C}$ to $50\mbox{${}^{\circ}$C}$—is shown by the violet bars. The upward arrows denote that a liquid ocean could exist above the corresponding temperatures. (Let us emphasize that temperatures of the Earth and Mars in this figure are normalized to the different values: $T_{\rm Ms0}=218\,K=-55\mbox{${}^{\circ}$C}$ and $T_{\rm Es0}=288\,K=15\mbox{${}^{\circ}$C}$, respectively [23].)
@@ 2.3 Analysis for the entire lifetime of the Earth and Mars
While the upper left panel in Fig. 1 represents a summary of observational data, three other panels demonstrate a confrontation of these data with theoretical curves (8) for three different cosmological models, characterized by the relative densities of the ‘dark energy’ ${\Omega}_{\Lambda}$ and non-relativistic (dust-like) matter ${\Omega}_{\rm D}$, both ‘dark’ and visible. (The Universe is always assumed to be spatially flat, so that ${\Omega}_{\Lambda}+{\Omega}_{\rm D}=1$.) The theoretical dependences $T_{\rm s}(t)/T_{\rm s0}$ are drawn by the black dotted curves, and the corresponding values of the contemporary Hubble parameter $H_{0}$ are written near each curve. For concineses, these values are normalized to $h=100$ km/s/Mpc.
In the case of commonly-accepted cosmological model with ${\Omega}_{\Lambda}=0.7$ and ${\Omega}_{\rm D}=0.3$ (bottom left panel), the best fit is achieved at $H_{0}=1.2{-}1.5~h=120{-}150$ km/s/Mpc, i.e., almost two times greater than the standard value of Hubble parameter at the intergalactic scales. The martian data (violet bars) suggest that such values of $H_{0}$ are consistent with the assumption of weakly-salted ocean, whose freezing point was above $10\mbox{${}^{\circ}$C}$. Significant outliers in the terrestrial data at the late times (within 1 billion years from today) are not surprising: this period involved an active biological evolution, resulting in the considerable variations of coefficients $A$ and $\alpha$ in our model (6), which were not taken into account.
@@
On the other hand, in the opposite case when the entire cosmological background is formed by the non-relativistic (dust-like) matter (${\Omega}_{\Lambda}=0$ and ${\Omega}_{\rm D}=1$, bottom right panel), the best-fitting Hubble parameter can be somewhat reduced: $H_{0}=1.0{-}1.2~h=100{-}120$ km/s/Mpc. Unfortunately, the corresponding theoretical curves become noticeably concaved and fit the experimental data worse at the intermediate times.